Part 1: Create an original example scenario to test your fellow classmates.

Your client wants to buy $500,000 The client has $50,000 to put down as a down payment, is approximately 32 years old, and is looking for a house with 4 bedrooms, 4.5 baths and close to the school. The client earns $5,200 per month and has no other debts and plans to stay at this location for 30 years. What kind of loan would you advise them to get and why?

Would the mortgage be conforming or nonconforming and why?
Answer #1

As he is planning to stay at the house for 30 years, he should buy a house for permanent residency.

Hence, he should go for a Home Loan.

He has $50000 and hence, he needs additional (500000-50000)=$450000 to own the house.

In USA, the average mortgage rate is 4.01% for 30 year loan. Hence,

Tenure(nper)=30*12=360, Loan amount(PV)=450000, rate=4.01%/12=$2150.96

E PV rate F G H 450000] 0.3342% 360 |=PMT(F2,F3,F1,0,0) PMT(rate, nper, pv, [fv], [type]) nper PMT

E PV rate nper PMT 450000 0.3342% 360 |($2,150.96)

Hence, As his earning is $5200/month and the monthly payment is $2150.96 , he can afford the loan.

And the mortgage would be conforming.